# Holistic Evaluation of Language Models

### v1

@@ Abstract
As language models (LMs) are increasingly adopted across domains, high-quality benchmarking frameworks that accurately estimate performance are essential for guiding deployment decisions. While frameworks such as Holistic Evaluation of Language Models (HELM) enable broad evaluation across tasks, they often rely on fixed prompts that fail to generalize across LMs, yielding unrepresentative performance estimates. Unless we estimate each LM’s ceiling (maximum achievable via changes to the prompt), we risk underestimating performance. Declarative prompting frameworks, such as DSPy, offer a scalable alternative to manual prompt engineering by crafting structured prompts that can be optimized per task. However, such frameworks have not been systematically evaluated across established benchmarks. We present a reproducible DSPy+HELM framework that introduces structured prompting methods which elicit reasoning, enabling more accurate LM benchmarking. Using four prompting methods, we evaluate four frontier LMs across seven benchmarks (general/medical domain) against existing HELM baseline scores. We find that without structured prompting: (i) HELM underestimates LM performance (by 4% average), (ii) performance estimates vary more across benchmarks (++2% standard deviation), (iii) performance gaps are misrepresented (leaderboard rankings flip on $3/7$ benchmarks), and (iv) introducing reasoning (chain-of-thought) reduces LM sensitivity to prompt design (smaller performance $\Delta$ across prompting methods). To our knowledge, this is the first large-scale benchmarking study to empirically characterize LM behavior across benchmarks and prompting methods, showing that scalable performance ceiling estimation enables more decision-useful benchmarks. We open-source (i) DSPy+HELM Integration [^1] DSPy+HELM Integration: https://github.com/stanford-crfm/helm/pull/3893 and (ii) Prompt Optimization Pipeline [^2] Prompt Optimization Pipeline: https://github.com/StanfordMIMI/dspy-helm.
@@ 1 Introduction
While benchmarking frameworks such as Holistic Evaluation of Language Models (HELM) (liang2022holistic; bedi2025medhelm) enable holistic evaluation via a comprehensive suite covering diverse tasks, public leaderboards typically evaluate multiple LMs under a fixed prompt per benchmark. However, fixed prompts rarely generalize well across LMs, leading to unrepresentative performance estimates that obscure underlying strengths and weaknesses of LMs. Hence, broader LM adoption necessitates scalable estimation of performance ceilings (i.e., the maximum achievable via prompt-only changes), thereby allowing practitioners to weigh cost–benefit tradeoffs and choose the right model for each downstream task.
@@
However, declarative prompting frameworks such as DSPy have not been systematically evaluated across established benchmarks, and there is limited guidance on how structured prompting compares with existing baselines. To address this gap, we integrate DSPy with HELM (Figure 1) and present:
@@
    An evaluation of prompting methods (Zero-Shot, Bootstrap Few-Shot with Random Search, MIPROv2) against HELM’s baseline across four LMs and seven HELM benchmarks that span general and medical domains (reasoning, knowledge QA, problem-solving, error-classification), where each prompting method leverages a distinct mechanism for refining prompts to estimate LM performance ceilings.
@@ 2.3 Experimental Setup
Implementation details. We evaluate four frontier LMs (Table 1) to study whether prompting meaningfully impacts the performance of even top-tier LMs. We seed each DSPy program with the baseline instruction for the corresponding benchmark from HELM’s prompt. For BFRS and MIPROv2 optimizers, we follow DSPy’s data separation: the demonstration pool is bootstrapped exclusively from the training split, while candidate prompts are evaluated on a disjoint held-out validation split from the original training partition; neither optimizer ever sees the HELM leaderboard test set. Each benchmark’s loader creates a fixed train/val partition (default 90/10 with the same seed), and we cap both bootstrapped and labeled demonstrations at $K\leq 3$ per module. All final scoring is performed via HELM, so outputs are judged identically regardless of how they were produced. All results reflect single, deterministic runs (temperature = 0), matching HELM’s experimental setup. For HELM baselines, we report HELM’s public leaderboard scores when the setup matches ours: (i) identical LM API version, (ii) zero-shot prompting, and (iii) no CoT reasoning. For benchmarks where the leaderboard setup does not match, we reproduce them with single, deterministic runs.
Metric calculation. To summarize overall gains, we take the mean of the three structured prompting methods (Zero-Shot CoT, BFRS, MIPROv2); for each LM, we first macro-average across benchmarks, and then average the $\Delta$ (absolute % change over baseline) across LMs. The change in variability is reported analogously using the per-model across-benchmark standard deviation, $\sigma$. For each model and prompting method, $\sigma$ is computed over per-benchmark scores (equal weight per benchmark).
@@ 3.1 Impact of Structured Prompting on HELM Leaderboard
Improved performance over HELM baseline. Across LMs and benchmarks, structured prompting methods (Zero-Shot CoT, BFRS, MIPROv2) consistently improve over the HELM baseline (Table 3). On average, LMs gain ++4% in absolute accuracy. Non-reasoning models benefit most (++5%), while the reasoning-optimized o3 Mini sees smaller but consistent gains (++2%).
Flipped leaderboard rankings. Taking ceiling performance into account alters three leaderboard rankings. On MMLU-Pro (Table 4), baseline o3 Mini > Claude 3.7 Sonnet (77.1% vs. 76.3%) reverses to Claude 3.7 Sonnet > o3 Mini (80.6% vs. 78.4%). On GSM8K, GPT 4o overtakes Gemini 2.0 Flash, shifting from (81.1% vs. 84.0%) to (90.7% vs. 84.2%). Finally, on MedCalc-Bench (Table 5), baseline o3 Mini > Claude 3.7 Sonnet (34.0% vs. 21.0%) becomes Claude 3.7 Sonnet > o3 Mini (35.3% vs. 34.7%), highlighting how prompting can act as a confounder in leaderboard rankings.
Altered inter-model performance gaps. When evaluated at ceiling performance (best score across prompting methods), models can either narrow or widen their relative performance gaps, providing a more accurate view of true capability differences. Averaging across benchmarks, the gap between the top two models (o3 Mini and Claude 3.7 Sonnet) shrinks from 6% at baseline (70.9 vs. 64.8) to 3% (73.2 vs. 69.8). However, this trend is not uniform: on GPQA, the gap widens substantially, from 0.6% at baseline (57.6 vs. 57.0) to 4.3% at ceiling (68.4 vs. 64.1).
@@
Ranking stability analysis. To assess how structured prompting shifts relative model rankings, we compute mean ranks (1 = best, 4 = worst) across all seven benchmarks. Under structured prompting, the ranking spread compresses: o3 Mini remains the top model but becomes less dominant (1.29 $\to$ 1.57), Claude 3.7 Sonnet improves and moves closer to the top (2.29 $\to$ 2.00), GPT 4o also improves modestly (3.14 $\to$ 3.00), while Gemini 2.0 Flash slightly declines (3.29 $\to$ 3.43). Rank standard deviation ($\sigma$) shows a similar but more moderate pattern: Claude 3.7 Sonnet (0.95 $\to$ 1.15) and GPT 4o (0.90 $\to$) 1.00) become slightly more volatile, while Gemini 2.0 Flash stabilizes (0.76 $\to$ 0.53), and o3 Mini remains roughly unchanged (0.76 $\to$ 0.79). While effects are modest (only 3 instances of ranking flips across leaderboards), the results demonstrate that benchmark rankings are not invariant to prompt design.
3.2 Chain-of-Thought vs Automatic Prompt Optimization
@@
Performance gains primarily attributed to CoT. We study the impact of each prompting method on the leaderboard by averaging results across LMs and benchmarks. Moving from HELM’s baseline to Zero-Shot Predict yields minimal improvement (64.6% $\to$ 64.9%). In contrast, introducing CoT reasoning and moving from Zero-Shot Predict to Zero-Shot CoT results in substantial gains (64.9% $\to$ 68.5%). Interestingly, moving from Zero-Shot CoT to more sophisticated optimizers, such as BFRS and MIPROv2, does not lead to a meaningful additional improvement (68.5% $\to$ 68.6%), indicating that once CoT is introduced, LMs become relatively insensitive to further optimization.
Why CoT reduces sensitivity to prompt design? We provide empirical evidence that reasoning LMs such as o3 Mini, or non-reasoning LMs with prompt-induced reasoning (CoT), are relatively insensitive to prompt design (meincke2025prompting). Now, we provide deeper insight into this phenomenon. With CoT, the model marginalizes over many reasoning paths; prompt optimization mostly reweights sampled paths. A model with parameters $\theta$ reads an input $x$ (a benchmark item) under a prompt $p$, samples a reasoning path $\tau$ (chain-of-thought), and produces a final answer $y$; we compare this to an alternative prompt $p^{\prime}$ (a structured prompt variant). CoT factorizes the joint as: (1) $P_{\theta}(\tau,y\mid x,p)=P_{\theta}(\tau\mid x,p)\,P_{\theta}(y\mid x,\tau,p)$ and predicts by marginalizing over paths (self-consistency): (2) $P_{\theta}(y\mid x,p)=\sum_{\tau}P_{\theta}(\tau\mid x,p)\,P_{\theta}(y\mid x,\tau,p)$. Because (2) is a marginalization of (1), the data-processing inequality implies that the change in the answer distribution between prompts $p$ and $p^{\prime}$ is upper-bounded by the change in the path distribution: $\big\|P_{\theta}(y\mid x,p)-P_{\theta}(y\mid x,p^{\prime})\big\|_{TV}\;\leq\;\big\|P_{\theta}(\tau\mid x,p)-P_{\theta}(\tau\mid x,p^{\prime})\big\|_{TV}\;\leq\;\sqrt{\tfrac{1}{2}\,D_{KL}\!\big(P_{\theta}(\tau\mid x,p)\,\|\,P_{\theta}(\tau\mid x,p^{\prime})\big)}$ where $\|\cdot\|_{\mathrm{TV}}$ is total-variation distance and $\mathrm{KL}(\cdot\|\cdot)$ is Kullback–Leibler divergence (applying the data-processing inequality followed by Pinsker’s inequality). We define the decision margin at item $x$ under prompt $p$ as $m(x;p)=P_{\theta}(y^{\star}\mid x,p)-\max_{y\neq y^{\star}}P_{\theta}(y\mid x,p),\quad y^{\star}=\arg\max_{y}P_{\theta}(y\mid x,p).$ If the left-hand side of the previous bound is $\tfrac{1}{2}\,m(x;p)$, then the prediction is invariant: $\arg\max_{y}P_{\theta}(y\mid x,p^{\prime})\;=\;\arg\max_{y}P_{\theta}(y\mid x,p).$ Since CoT enlarges $m(x;p)$ by averaging diverse valid paths, typical prompt optimization, which mainly reweights $P_{\theta}(\tau\mid x,p)$, cannot flip decisions except on near-tied items, making reasoning models largely prompt-insensitive. If $\mathrm{KL}\!\big(P_{\theta}(\tau\mid x,p)\,\|\,P_{\theta}(\tau\mid x,p^{\prime})\big)\leq\kappa$ and $m(x;p)\geq 2\varepsilon$, then predictions are invariant under $p\!\to\!p^{\prime}$ whenever $\sqrt{\kappa/2}<\varepsilon$.
@@ 4 Related Work
Holistic benchmarking. The General Language Understanding Evaluation (GLUE) (wang2018glue) benchmark was one of the first multi-task evaluation frameworks, aggregating nine distinct language understanding tasks. Benchmarks of increasing scale followed: (i) Measuring Massive Multitask Language Understanding (MMLU) (hendrycks2021measuringmassivemultitasklanguage), including 57 tasks spanning STEM, humanities, social sciences, and (ii) Beyond the Imitation Game (BIG-Bench) (srivastava2023imitationgamequantifyingextrapolating), with 204 diverse tasks. The HELM framework is an established standard, designed for transparent, reproducible, and multi-metric evaluation of model capabilities (liang2022holistic). However, these benchmarks are typically evaluated using static prompts. liang2022holistic note they opt for simple, generic prompts to orient development "towards generic language interfaces" that do not require "model-specific incantations". This reliance on fixed prompts, however, risks the underestimation of the true capabilities of LMs. srivastava2023imitationgamequantifyingextrapolating conclude that standard few-shot prompting "substantially underestimates the best performance and capabilities of language models" when compared to methods that elicit reasoning (suzgun2022challengingbigbenchtaskschainofthought).
Prompting methods. The discovery of in-context learning (brown2020languagemodels), where models learn from n-shot demonstrations, and the breakthrough of chain-of-thought (CoT) prompting (wei2022chainofthought) established the important role of prompt design in model performance. Complex, manually-composed strategies like Medprompt (nori2023medprompt), which combine few-shot selection, CoT, and ensembling, demonstrate that an LM’s performance ceiling often lies higher than with the use of static prompts. Because manual prompt engineering is impractical for systematically finding this ceiling, researchers often frame prompt design as a formal "optimization problem", leading to the field of APO. Early APO methods include generation-and-selection, such as Automatic Prompt Engineer (APE) (zhou2023largelanguagemodelshumanlevel), which uses an LM to propose candidate instructions and a separate scoring function to select the best one. Subsequent systems expanded this search paradigm (wang2023promptagent; yang2023large; singla2024dynamic). These methods often outperform zero-shot or manually engineered prompts on a variety of general tasks. In the LM-as-Optimizer paradigm, an LM is instructed to iteratively refine prompts by showing it a trajectory of previously evaluated candidates and their scores. Other approaches have employed evolutionary search, like Promptbreeder (fernando2024promptbreeder), which treats prompts as "genes" and evolves a population of instructions over generations using a LM to perform mutation. The DSPy framework (khattab2023dspy) generalizes these methods, providing a programming model that compiles declarative, multi-stage pipelines.

### v2

# Structured Prompting Enables More Robust Evaluation of Language Models

@@ Abstract
As language models (LMs) are increasingly adopted across domains, high-quality benchmarking frameworks that accurately estimate performance are essential for guiding deployment decisions. While frameworks such as Holistic Evaluation of Language Models (HELM) enable broad evaluation across tasks, they often rely on fixed prompts that fail to generalize across LMs, yielding unrepresentative performance estimates. Unless we approximate each LM’s ceiling (maximum achievable via changes to the prompt), we risk underestimating performance. Declarative prompting frameworks, such as DSPy, offer a scalable alternative to manual prompt engineering by crafting structured prompts that can be optimized per task. However, such frameworks have not been systematically evaluated across established benchmarks. We present a reproducible DSPy+HELM framework that introduces structured prompting methods which elicit reasoning, enabling more accurate LM benchmarking. Using four prompting methods, we evaluate four frontier LMs across seven benchmarks (general/medical domain) against existing HELM baseline scores. We find that without structured prompting: (i) HELM underestimates LM performance (by 4% average), (ii) performance estimates vary more across benchmarks (++2% standard deviation), (iii) performance gaps are misrepresented (leaderboard rankings flip on $3/7$ benchmarks), and (iv) introducing reasoning (chain-of-thought) reduces LM sensitivity to prompt design (smaller performance $\Delta$ across prompting methods). To our knowledge, this is the first benchmarking study to systematically integrate structured prompting into an established evaluation framework, demonstrating how scalable performance-ceiling approximation yields more robust, decision-useful benchmarks. We open-source (i) DSPy+HELM Integration [^1] DSPy+HELM Integration: https://github.com/stanford-crfm/helm/pull/3893 and (ii) Prompt Optimization Pipeline [^2] Prompt Optimization Pipeline: https://github.com/StanfordMIMI/dspy-helm.
@@ 1 Introduction
While benchmarking frameworks such as Holistic Evaluation of Language Models (HELM) (liang2022holistic; bedi2025medhelm) enable holistic evaluation via a comprehensive suite covering diverse tasks, public leaderboards typically evaluate multiple LMs under a fixed prompt per benchmark. However, fixed prompts rarely generalize well across LMs, leading to unrepresentative performance estimates that obscure underlying strengths and weaknesses of LMs. Hence, broader LM adoption necessitates scalable approximation of performance ceilings (i.e., the maximum achievable via prompt-only changes), thereby allowing practitioners to weigh cost–benefit tradeoffs and choose the right model for each downstream task.
@@
However, despite the growing use of structured prompting, we lack a systematic evaluation of how these approaches affect benchmark robustness and performance estimates across established evaluation suites. We use DSPy as an instantiation of structured prompting and integrate it with HELM (Figure 1), presenting:
@@
    An evaluation of prompting methods (Zero-Shot, Bootstrap Few-Shot with Random Search, MIPROv2) against HELM’s baseline across four LMs and seven HELM benchmarks that span general and medical domains (reasoning, knowledge QA, problem-solving, error-classification), where each prompting method leverages a distinct mechanism for refining prompts to approximate LM performance ceilings.
@@ 2.3 Experimental Setup
Implementation details. We evaluate four frontier LMs (Table 1). We initialize each DSPy program with HELM’s baseline instruction for comparability. DSPy then applies its own standardized prompting modules, treating the full HELM prompt as input. For BFRS and MIPROv2 optimizers, we follow DSPy’s data separation: the demonstration pool is bootstrapped exclusively from the training split, while candidate prompts are evaluated on a disjoint held-out validation split from the original training partition; neither optimizer ever sees the HELM leaderboard test set. Each benchmark’s loader creates a fixed train/val partition (default 90/10 with the same seed), and we cap both bootstrapped and labeled demonstrations at $K\leq 3$ per module. All final scoring is performed via HELM, so outputs are judged identically regardless of how they were produced. All results reflect single, deterministic runs (temperature = 0), matching HELM’s experimental setup. For HELM baselines, we report HELM’s public leaderboard scores when the setup matches ours: (i) identical LM API version, (ii) zero-shot prompting, and (iii) no CoT reasoning. For benchmarks where the leaderboard setup does not match, we reproduce them with single, deterministic runs.
Metric calculation. To summarize gains, we take the mean of the three structured prompting methods (Zero-Shot CoT, BFRS, MIPROv2); for each LM, we first macro-average across benchmarks, and then average the $\Delta$ (absolute % change over baseline) across LMs. The change in variability ($\sigma$) is reported analogously.
@@ 3.1 Impact of Structured Prompting on HELM Leaderboard
Improved performance over HELM baseline. Structured prompting methods (Zero-Shot CoT, BFRS, MIPROv2) consistently improve over the HELM baseline (Table 3). On average, LMs gain ++4% in absolute accuracy. Non-reasoning models benefit most (++5%), while o3 Mini sees smaller but consistent gains (++2%).
Flipped leaderboard rankings. At ceiling, three leaderboard rankings flip. On MMLU-Pro (Table 4), baseline o3 Mini > Claude 3.7 Sonnet (77.1% vs. 76.3%) reverses to Claude 3.7 Sonnet > o3 Mini (80.6% vs. 78.4%). On GSM8K, GPT 4o overtakes Gemini 2.0 Flash, shifting from (81.1% vs. 84.0%) to (90.7% vs. 84.2%). On MedCalc-Bench (Table 5), baseline o3 Mini > Claude 3.7 Sonnet (34.0% vs. 21.0%) becomes Claude 3.7 Sonnet > o3 Mini (35.3% vs. 34.7%), highlighting how prompt choice can moderate rankings.
Altered inter-model performance gaps. When evaluated at ceiling performance, models can either narrow or widen their relative performance gaps, providing a more accurate view of true capability differences. Averaging across benchmarks, the gap between the top two models (o3 Mini and Claude 3.7 Sonnet) shrinks from 6% at baseline (70.9 vs. 64.8) to 3% (73.2 vs. 69.8). However, this trend is not uniform: on GPQA, the gap widens substantially, from 0.6% at baseline (57.6 vs. 57.0) to 4.3% at ceiling (68.4 vs. 64.1).
@@
Ranking stability analysis. To assess how structured prompting shifts relative rankings, we compute mean ranks (1 = best, 4 = worst) across all seven benchmarks. Under structured prompting, the ranking spread compresses: o3 Mini remains the top model but becomes less dominant (1.29 $\to$ 1.57), Claude 3.7 Sonnet improves and moves closer to the top (2.29 $\to$ 2.00), GPT 4o also improves modestly (3.14 $\to$ 3.00), while Gemini 2.0 Flash slightly declines (3.29 $\to$ 3.43). Rank standard deviation ($\sigma$) shows a similar but more moderate pattern: Claude 3.7 Sonnet (0.95 $\to$ 1.15) and GPT 4o (0.90 $\to$) 1.00) become slightly more volatile, while Gemini 2.0 Flash stabilizes (0.76 $\to$ 0.53), and o3 Mini remains roughly unchanged (0.76 $\to$ 0.79). Overall, the results demonstrate that leaderboard rankings are not invariant to prompt design.
@@
CoT reduces sensitivity to prompt design. We study the impact of each prompting method on the leaderboard by averaging results across LMs and benchmarks. Moving from HELM’s baseline to Zero-Shot Predict yields minimal improvement (64.6% $\to$ 64.9%). In contrast, introducing CoT reasoning and moving from Zero-Shot Predict to Zero-Shot CoT results in substantial gains (64.9% $\to$ 68.5%). Interestingly, moving from Zero-Shot CoT to more sophisticated optimizers, such as BFRS and MIPROv2, does not lead to a meaningful additional improvement (68.5% $\to$ 68.6%), indicating that once CoT is introduced, LMs become less sensitive to further optimization.
3.2 Theoretical Insight: "Why CoT Reduces Sensitivity to Prompt Design"
We first formalize the effect of CoT on prompt sensitivity. Consider a LM with parameters $\theta$, input $x$, and two prompts $p$ and $p^{\prime}$ that share the same CoT interface but differ in instructions and/or demonstrations [^3] Throughout, we fix the decoding temperature and sampling strategy, so that changing $p$ only affects the textual prefix.. Under prompt $p$, the model samples a full reasoning path $\tau$ (CoT) and then produces a final answer $y$; i.e.,
$P_{\theta}(\tau,y\mid x,p)=P_{\theta}(\tau\mid x,p)\,P_{\theta}(y\mid x,\tau,p).$ (2)
The predictive answer distribution under $p$ is obtained by marginalizing over reasoning paths (self-consistency):
$P_{\theta}(y\mid x,p)=\sum_{\tau}P_{\theta}(\tau\mid x,p)\,P_{\theta}(y\mid x,\tau,p).$ (3)
Once a full reasoning path $\tau$ has been generated, the residual dependence of $y$ on the prompt is negligible:
$P_{\theta}(y\mid x,\tau,p)\approx P_{\theta}(y\mid x,\tau,p^{\prime})\approx P_{\theta}(y\mid x,\tau).$ (4)
Because all structured prompt variants instruct the LM to output a reasoning trace, once $\tau$ is fixed, small changes in the instructions/demonstrations do not systematically change the conditional distribution over $y$:
$p\to\tau\to y\quad\text{forms a Markov chain given $x$},$ (5)
i.e., $y\perp p\mid(x,\tau)$. The answer distribution $P_{\theta}(y\mid x,p)$ is obtained by passing the path distribution $P_{\theta}(\tau\mid x,p)$ through a fixed channel $P_{\theta}(y\mid x,\tau)$. Let $\|\cdot\|_{\mathrm{TV}}$ denote total variation distance and $D_{\mathrm{KL}}(\cdot\|\cdot)$ Kullback–Leibler divergence. Because $P_{\theta}(y\mid x,p)$ is the image of $P_{\theta}(\tau\mid x,p)$ under $\tau\mapsto y$, the data-processing inequality yields
$\big\|P_{\theta}(y\mid x,p)-P_{\theta}(y\mid x,p^{\prime})\big\|_{\mathrm{TV}}\;\leq\;\big\|P_{\theta}(\tau\mid x,p)-P_{\theta}(\tau\mid x,p^{\prime})\big\|_{\mathrm{TV}}.$ (6)
Applying Pinsker’s inequality to the right-hand side gives
$\big\|P_{\theta}(y\mid x,p)-P_{\theta}(y\mid x,p^{\prime})\big\|_{\mathrm{TV}}\;\leq\;\sqrt{\tfrac{1}{2}\,D_{\mathrm{KL}}\!\big(P_{\theta}(\tau\mid x,p)\,\|\,P_{\theta}(\tau\mid x,p^{\prime})\big)}.$ (7)
Thus, the extent to which the answer distribution can change under prompt perturbations is upper-bounded by how much the CoT path distribution changes. For a given prompt $p$ and item $x$, define the decision margin
$m(x;p)\;=\;P_{\theta}(y^{\star}\mid x,p)-\max_{y\neq y^{\star}}P_{\theta}(y\mid x,p),\qquad y^{\star}=\arg\max_{y}P_{\theta}(y\mid x,p).$ (8)
We now state a pointwise decision-stability result. Fix $x$ and prompts $p,p^{\prime}$. If
$\big\|P_{\theta}(y\mid x,p)-P_{\theta}(y\mid x,p^{\prime})\big\|_{\mathrm{TV}}\;<\;\tfrac{1}{2}\,m(x;p),$ (9)
then the prediction is invariant:
$\arg\max_{y}P_{\theta}(y\mid x,p^{\prime})\;=\;\arg\max_{y}P_{\theta}(y\mid x,p).$ (10)
Moreover, a sufficient condition is
[…]
Condition 9 implies that the probability mass on the top-class $y^{\star}$ cannot be reduced by more than $\tfrac{1}{2}m(x;p)$, while the mass on any competitor cannot be increased by more than the same amount. Hence no competitor can overtake $y^{\star}$, giving 10. Inequality 11 combines the TV bound in 7 with the margin condition 9.
In CoT decoding, $P_{\theta}(y\mid x,p)$ can be viewed as a marginalization over possible reasoning paths, which typically enlarges the margin $m(x;p)$ compared to direct (non-CoT) decoding. At the same time, structured prompting methods mainly alter instructions and few-shot examples while preserving the CoT interface, so they primarily act by reweighting $P_{\theta}(\tau\mid x,p)$ rather than changing the conditional channel $P_{\theta}(y\mid x,\tau)$. Once CoT is enabled, the effective KL divergence between path distributions under different structured prompts is small enough that equation 11 holds for most items, and further optimization rarely flips decisions except on near-tied examples. Empirically, this is reflected in our results: moving from non-CoT to Zero-Shot CoT yields majority gains, while more aggressive optimizers (BFRS, MIPROv2) produce marginal improvements.
@@ 4 Related Work
Holistic benchmarking. The General Language Understanding Evaluation (GLUE) (wang2018glue) benchmark was one of the first multi-task evaluation frameworks, aggregating nine distinct language understanding tasks. Benchmarks of increasing scale followed: (i) Measuring Massive Multitask Language Understanding (MMLU) (hendrycks2021measuringmassivemultitasklanguage), including 57 tasks spanning STEM, humanities, social sciences, and (ii) Beyond the Imitation Game (BIG-Bench) (srivastava2023imitationgamequantifyingextrapolating), with 204 diverse tasks. The HELM framework is an established standard, designed for transparent, reproducible, and multi-metric evaluation of model capabilities (liang2022holistic). However, these benchmarks are typically evaluated using static prompts. liang2022holistic note they opt for simple, generic prompts to orient development "towards generic language interfaces" that do not require "model-specific incantations". This reliance on fixed prompts, however, risks the underestimation of the true capabilities of LMs. srivastava2023imitationgamequantifyingextrapolating; suzgun2022challengingbigbenchtaskschainofthought conclude that standard few-shot prompting substantially underestimates the capabilities of LMs.
Prompting methods. The discovery of in-context learning (brown2020languagemodels), where models learn from n-shot demonstrations, and the breakthrough of chain-of-thought (CoT) prompting (wei2022chainofthought) established the important role of prompt design in model performance. Complex, manually-composed strategies like Medprompt (nori2023medprompt), which combine few-shot selection, CoT, and ensembling, demonstrate that a LM’s performance ceiling often lies higher than with the use of static prompts. Because manual prompt engineering is impractical for systematically approximating this ceiling, researchers often frame prompt design as a formal "optimization problem", leading to the field of APO. Early APO methods include generation-and-selection, such as Automatic Prompt Engineer (APE) (zhou2023largelanguagemodelshumanlevel), which uses an LM to propose candidate instructions and a separate scoring function to select the best one. Subsequent systems expanded this search paradigm (wang2023promptagent; yang2023large; singla2024dynamic). These methods often outperform zero-shot or manually engineered prompts on a variety of general tasks. In the LM-as-Optimizer paradigm, an LM is instructed to iteratively refine prompts by showing it a trajectory of previously evaluated candidates and their scores. Other approaches have employed evolutionary search, like Promptbreeder (fernando2024promptbreeder), which treats prompts as "genes" and evolves a population of instructions over generations using a LM to perform mutation. The DSPy framework (khattab2023dspy) generalizes these methods, providing a programming model that compiles declarative, multi-stage pipelines.